# High-dimensional dynamics in low-dimensional networks

### v1

@@ Acknowledgments
Figure S.11: Response properties of a weakly low-rank network with $c>0$. Same as Figure 1 except $c=0.75$.
Figure S.12: Response properties of a weakly low-rank network with $c<0$. Same as Figure 1 except $c=-0.75$.
S.2 Mathematical analysis of low-rank networks with external perturbations
In this section we present a mathematical analysis of low-rank suppression and high-dimensional dynamics in networks with effective low-rank connectivity and external perturbations. Our analysis applies to a broad class of dynamics. We begin by reviewing the precise assumptions made in our analysis.
S.2.1 Assumptions on the dynamics and definition of the effective connectivity matrix.
We start by assuming dynamics of the form
$\tau\frac{d{\bf{z}}}{dt}={\bf{F}}({\bf{z}},{\bf{x}})$ (7)
for some smooth ${\bf{F}}:{\mathbb{R}}^{N}\times{\mathbb{R}}^{N}\to{\mathbb{R}}^{N}$ where $\tau>0$ sets the timescale of the dynamics. Here, ${\bf{z}}(t)$ is an $N\times 1$ vector quantifying the state of $N$ interacting nodes and ${\bf{x}}(t)$ is an $N\times 1$ vector of external forces, inputs, and/or noise. At the end of Section S.2.3, we show how our approach can be generalized to a larger class of causal, time invariant systems, but the dynamics in Eq. (7) are sufficient for all of the examples studied in this manuscript, and the majority of models studied in the literature on low-rank networks, so we focus on this class of dynamics for simplicity.
In general, it is not immediately clear how to define a network of interactions, $W$, from dynamics as general as those in Eq. (7). For example, in nonlinear systems, the effective connectivity between nodes can change with the state of the system. To circumvent this problem, we define the effective connectivity matrix, $W$, through a linearization around a stable equilibrium.
Specifically, we assume that there exists a time-constant forcing, ${\bf{x}}(t)={\bf{x}}_{0}$, that produces a stable steady state solution, ${\bf{z}}_{0}$. In other words,
${\bf{F}}({\bf{z}}_{0},{\bf{x}}_{0})={\bf{0}}.$
$J_{{\bf{z}}}=\partial_{{\bf{z}}}F({\bf{z}}_{0},{\bf{x}}_{0})$
has eigenvalues with strictly negative real part. We then consider a small perturbation around this fixed point driven by a perturbation to the forcing term,
${\bf{x}}_{p}(t)={\bf{x}}_{0}+\epsilon{\bf{x}}(t).$
The response, ${\bf{z}}_{p}(t)$, of the network to the perturbed forcing term, ${\bf{x}}_{p}(t)$, can be written to linear order in $\epsilon$ as
${\bf{z}}_{p}(t)={\bf{z}}_{0}+\epsilon{\bf{z}}+\mathcal{O}(\epsilon^{2}).$ (8)
The perturbation, ${\bf{z}}(t)$, obeys the linearized equation
$\tau\frac{d{\bf{z}}}{dt}=-{\bf{z}}+W{\bf{z}}+J_{{\bf{x}}}{\bf{x}}(t)$ (9)
where $J_{{\bf{x}}}=\partial_{{\bf{x}}}{\bf{F}}({\bf{z}}_{0},{\bf{x}}_{0})$ is the Jacobian matrix of ${\bf{F}}({\bf{z}},{\bf{x}})$ with respect to ${\bf{x}}$ evaluated at the fixed point, and
$W=J_{{\bf{z}}}+I$ (10)
with $I$ the identity matrix and $J_{{\bf{z}}}=\partial_{{\bf{z}}}{\bf{F}}({\bf{z}}_{0},{\bf{x}}_{0})$ the Jacobian matrix of ${\bf{F}}({\bf{z}},{\bf{x}})$ with respect to ${\bf{z}}$ evaluated at the fixed point. We interpret $W$ in Eq. (10) as the effective connectivity matrix.
It may seem counterintuitive at first to define $W=J_{{\bf{z}}}+I$ instead of $W=J_{{\bf{z}}}$. To understand our choice, note that when we choose $W=J_{{\bf{z}}}+I$, the perturbed system ${\bf{z}}_{p}(t)$ decays to the unperturbed fixed point, ${\bf{z}}_{0}(t)$, whenever there is no perturbation or connectivity (${\bf{x}}(t)=0$ and $W=0$). Moreover, when there is a perturbation without connectivity ($W=0$ and ${\bf{x}}\neq 0$), the perturbation converges to ${\bf{z}}=J_{{\bf{x}}}{\bf{x}}$. Hence, $W$ represents interactions that determine the dynamics under perturbations away from the fixed point. In any case, our overall results are not sensitive to this choice.
One caveat to our approach for defining effective connectivity, $W$, from the dynamics in Eq. (7) is that the input perturbation, ${\bf{x}}(t)$, must be weak enough and/or the dynamics must be sufficiently linear for the linearized dynamics to be accurate, i.e., for the $\mathcal{O}(\epsilon^{2})$ term in Eq. (8) to be small. Moreover, the dynamics of the unperturbed system must be stable near the steady state. Examples with unstable, nonlinear dynamics are considered in Supplementary Figures S.7 and S.8.
Another caveat is that multiplication of ${\bf{x}}$ by $J_{{\bf{x}}}$ in Eq. (9) could suppress dimensionality, for example if $J_{{\bf{x}}}$ is low-rank or effectively low-rank. This would correspond to an effectively feedforward “read-in” mechanism that lowers dimensionality. Specifically, the “effective input” $J_{{\bf{x}}}{\bf{x}}$ would be low-dimensional and suppressed in some directions. For example, consider dynamics of the form
$\tau\frac{d{\bf{z}}}{dt}=-{\bf{z}}+W{\bf{z}}+W_{x}{\bf{x}}$ (11)
where $W_{x}$ is some matrix. In this model, we have $J_{{\bf{x}}}=W_{x}$. Therefore, if $W_{x}$ were low-rank then the effective input, ${\bf{x}}_{eff}=W_{x}{\bf{x}}$, would be low-dimensional even when the raw input, ${\bf{x}}$, were high-dimensional. This scenario would promote low-dimensional dynamics in ${\bf{z}}(t)$. An example of this form is considered in Supplementary Figure S.9. In many models (including Eqs. (1) and (3)), the dynamics take the form ${\bf{F}}({\bf{z}},{\bf{x}})={\bf{F}}({\bf{z}})+{\bf{x}}$ so that $J_{{\bf{x}}}=I$ and this is not an issue. More generally, we assume that $J_{{\bf{x}}}$ is full rank in all of our analysis.
Before continuing, we compare our approach for relating dynamics to connectivity with the approach taken in Thibeault et al. [1]. They define dynamics that explicitly take the connectivity matrix into account,
$\frac{d{\bf{z}}}{dt}$ $={\bf{G}}({\bf{z}},{\bf{y}})$
${\bf{y}}$ $=U{\bf{z}}$
where $U$ is interpreted as the connectivity matrix (they use “$W$”, but we switched to “$U$” to avoid confusion with our definition). This formulation represents a self-contained network in the absence of external perturbations or noise, but it can easily be extended to allow external perturbations by defining
$\frac{d{\bf{z}}}{dt}={\bf{G}}({\bf{z}},{\bf{y}},{\bf{x}}).$ (12)
However, Eq. (12) is not suitable for our purposes because $U$ does not necessarily capture all of the interactions between nodes. For example, the $j$th entry of ${\bf{G}}({\bf{z}},{\bf{y}},{\bf{x}})$ could depend explicitly (through its first argument) on the $k$th entry of ${\bf{z}}$ for some $j\neq k$. This would represent an interaction between nodes $j$ and $k$, which are not accounted for in $U$. For example, we could have $d{\bf{z}}_{j}/dt={{\bf{G}}}_{j}({\bf{z}},{\bf{y}},{\bf{x}})={\bf{z}}_{k}+{\bf{y}}_{j}={\bf{z}}_{k}+U_{ji}{\bf{z}}_{i}$ for three distinct indices $i,j,k$ where we assume that $U_{jk}=0$. In this case, the connectivity matrix, $U$, correctly captures the influence of node ${\bf{z}}_{i}$ on node ${\bf{z}}_{j}$ through the entry $U_{ji}$. However, node ${\bf{z}}_{j}$ is also influenced by node ${\bf{z}}_{k}$ and this interaction is not captured by the matrix $U$ (as we take $U_{jk}=0$) because it is instead captured by the explicit dependence of ${\bf{G}}$ on ${\bf{z}}_{k}$ (the first argument of ${\bf{G}}$). This could, of course, be overcome by incorporating the dependence on ${\bf{z}}_{k}$ into $U$ (i.e., by setting $U_{jk}=1$), but Eq. (12) makes no assumption that all such interactions are included into $U$. More importantly, it might not be possible to account for higher order, nonlinear interactions using the formulation in Eq. (12). For example, we could have $d{\bf{z}}_{j}/dt={\bf{G}}_{j}({\bf{z}},{\bf{y}},{\bf{x}})={\bf{y}}_{k}{\bf{z}}_{i}=U_{jk}{\bf{z}}_{k}{\bf{z}}_{i}$ for three distinct indices $i,j,k$. In this case, the interaction between ${\bf{z}}_{i}$ and ${\bf{z}}_{j}$ would not be accounted for in the matrix $U$, and it is not possible to account for both interactions ($i\to j$ and $k\to j$) in the matrix $U$ using the formulation in Eq. (12). Moreover, the strength of the interaction between ${\bf{z}}_{i}$ and ${\bf{z}}_{j}$ depends on the magnitude of ${\bf{z}}_{k}$, which can change with the state of the network. These matters are handled in the mathematical analysis of Thibeault et al. [1] by bounding norms related to the Jacobian matrix of ${\bf{G}}$ with respect to ${\bf{z}}$ and ${\bf{y}}$, and they also separately consider higher order interactions. Our approach avoids these bounds on the Jacobian matrices at the expense of relying on a linearization near a steady state. Under the formulation from Eq. (12), if there is a fixed point at ${\bf{z}}_{0}$ in response to static input ${\bf{x}}_{0}$, then our interpretation of the connectivity matrix is given by
$W=\partial_{{\bf{z}}}{\bf{G}}({\bf{z}}_{0},{\bf{y}}_{0},{\bf{x}}_{0})+\partial_{{\bf{y}}}{\bf{G}}({\bf{z}}_{0},{\bf{y}}_{0},{\bf{x}}_{0})U+I$
where ${\bf{y}}_{0}=U{\bf{z}}_{0}$ and the two partial derivatives denote Jacobian matrices with respect to the first and second arguments, respectively. Hence, the two approaches can easily be directly related.
In summary, we consider the dynamics defined by Eq. (7), linearized around a stable steady state, which ultimately gives rise to the definition of network dynamics in Eq. (9). This approach has some caveats, which are discussed above.
S.2.2 Assumptions on the network structure.
The network represents $N$ interacting nodes, the interactions are defined by an $N\times N$ connectivity matrix, $W$, and we consider asymptotics in the $N\to\infty$ limit. Here, we specify the assumptions made about the structure of $W$ in the large $N$ limit.
Letting $\sigma_{1},\ldots,\sigma_{N}$ be the singular values of $W$, we assume that $\sigma_{k}\gg 1$ for $k=1,\ldots,r$ (in the sense that $\sigma_{k}\to\infty$ as $N\to\infty$) and that $\sigma_{k}\leq\mathcal{O}(1)$ for $k=r+1,\ldots,N$ (in the sense that $\sigma_{k}$ has a finite limit as $N\to\infty$). We also assume that $r\sim\mathcal{O}(1)$ so that $r\ll N$. In summary, $W$ has a small number of large singular values and the remaining singular values are much smaller.
Under these assumptions, $W$ can be decomposed as [1]
$W=W_{0}+W_{1}$
where $W_{0}$ has rank $r$ and we write its singular value decomposition as
$W_{0}=U\Sigma V^{T}.$
Here, $U$ and $V$ are $N\times r$ orthonormal matrices and $\Sigma$ is an $r\times r$ diagonal matrix with diagonal entries, $\Sigma_{kk}=\sigma_{k}$. The left and right singular matrices, $U$ and $V$, can be written as
$U=\left[{\bf{u}}_{1}\;\;{\bf{u}}_{2}\;\cdots\;{\bf{u}}_{r}\right]$
and
$V=\left[{\bf{v}}_{1}\;\;{\bf{v}}_{2}\;\cdots\;{\bf{v}}_{r}\right]$
where their columns, ${\bf{u}}_{k}$ and ${\bf{v}}_{k}$, are the left and right singular vectors, which satisfy $\|{\bf{u}}_{k}\|=\|{\bf{v}}_{k}\|=1$ and ${\bf{u}}_{k}\cdot{\bf{u}}_{j}={\bf{v}}_{k}\cdot{\bf{v}}_{j}=0$ for $j\neq k$. We further assume that $W_{1}$ is a random matrix, which is almost surely full rank and is statistically independent from $W_{0}$. Finally, we assume that the eigenvalues and singular values of $W_{1}$ are smaller than $1$ in magnitude and bounded away from $1$ ($\sigma_{W_{1}},|\lambda_{W_{1}}|<1-\delta$ for some fixed $\delta>0$) with probability $1$.
S.2.3 Analysis of low-rank suppression.
We now analyze the conditions under which low-rank suppression occurs. Our goal here is to understand the conditions under which the dynamics produce low-rank suppression like that observed in Figure 1c. In the next section, we will derive conditions on high-dimensional dynamics in the presence of stochastic perturbations such as those considered in Figure 1d–g.
We begin by considering the case in which ${\bf{x}}(t)$ is static or slow, so we can use a steady state or quasi-steady state approximation. Later in this section, we consider low-rank suppression with transient and fast perturbations.
If ${\bf{x}}(t)={\bf{x}}$ is static (such as during the duration of each stimulus in Figure 1c), then the network response converges to solutions of the perturbed steady state, which from Eq. (9) satisfies
${\bf{z}}=W{\bf{z}}+J_{{\bf{x}}}{\bf{x}}$
and therefore
${\bf{z}}=[I-W]^{-1}J_{{\bf{x}}}{\bf{x}}.$ (13)
If ${\bf{x}}(t)$ is not static, but changes much more slowly than $\tau$, then ${\bf{z}}(t)$ closely tracks the quasi-steady state given by Eq. (13).
Low-rank suppression, like that observed in Figure 1c, occurs whenever there is an input, ${\bf{x}}$, that is strongly suppressed by the recurrent dynamics in comparison to a random input (as in Figure 1c). In other words, we can say that low-rank suppression occurs whenever there is a raw input, ${\bf{x}}$, such that $\|{\bf{x}}\|=\mathcal{O}(1)$ and
$\|{\bf{z}}\|=\|[I-W]^{-1}J_{{\bf{x}}}{\bf{x}}\|\ll\|{\bf{z}}_{rand}\|=\|[I-W]^{-1}{\bf{u}}_{rand}\|$
with high probability whenever ${\bf{u}}_{rand}$ is a random unit vector, independent from $W_{0}$, $W_{1}$, and $J_{{\bf{x}}}$. To simplify this statement, first note that $\|[I-W]^{-1}{\bf{u}}_{rand}\|=\mathcal{O}(1)$ whenever ${\bf{u}}_{rand}$ is independent from $W_{1}$. Moreover, since we assume that $J_{{\bf{x}}}$ is full rank, we can write the condition in terms of the effective input,
${\bf{x}}_{eff}=J_{{\bf{x}}}{\bf{x}}$
instead of the raw input, ${\bf{x}}$.
In particular, low-rank suppression occurs whenever there is a vector ${\bf{x}}_{eff}$ such that
$\|{\bf{z}}\|=\|[I-W]^{-1}{\bf{x}}_{eff}\|\ll\|{\bf{x}}_{eff}\|.$ (14)
We take this property to be the definition of low-rank suppression.
For the dynamics that we use in our simulations, note that $J_{{\bf{x}}}=I$ so ${\bf{x}}_{eff}={\bf{x}}$. We discuss some caveats of the distinction between ${\bf{x}}$ and ${\bf{x}}_{eff}$ at the end of this section. We now prove that low-rank suppression arises under the assumptions from Sections S.2.2S.2.3.
Claim 1.
Under the model and assumptions of Sections S.2.2S.2.3, the network exhibits low-rank suppression.
Proof.
Let $k\in\{1,\ldots,r\}$ and take
${\bf{x}}_{eff}={\bf{u}}_{k}+\frac{W_{1}{\bf{v}}_{k}-{\bf{v}}_{k}}{\sigma_{k}}.$ (15)
We claim that the response to this input is given by
${\bf{z}}=-\frac{{\bf{v}}_{k}}{\sigma_{k}}.$
To see this, we only need to show that $[I-W]{\bf{z}}={\bf{x}}$. We have
${\bf{z}}$ $=\frac{W_{0}{\bf{v}}_{k}+W_{1}{\bf{v}}_{k}-{\bf{v}}_{k}}{\sigma_{k}}$
$=\frac{U\Sigma V^{T}{\bf{v}}_{k}+W_{1}{\bf{v}}_{k}-{\bf{v}}_{k}}{\sigma_{k}}$
$={\bf{u}}_{k}+\frac{W_{1}{\bf{v}}_{k}-{\bf{v}}_{k}}{\sigma_{k}}$
$={\bf{x}}_{eff}.$
Now note that $\|{\bf{x}}_{eff}\|=1+o(1)$ and $\|{\bf{z}}\|\ll 1$ because $\|{\bf{u}}_{k}\|=\|{\bf{v}}_{k}\|=1$, $\|W_{1}\|_{2}=\mathcal{O}(1)$, and $\sigma_{k}\gg 1$. Therefore $\|{\bf{z}}\|\ll\|{\bf{x}}_{eff}\|$. ∎
Before continuing, we will make a few observations about this claim and its proof. First note that the definition of low-rank suppression in Eq. (14) is equivalent to the statement that the matrix
$A=[I-W]^{-1}$
has at least one asymptotically small singular value,
$\sigma_{A}\ll 1.$
Singular values commute with matrix inverses, so the singular values of $A$ are given by
$\sigma_{A}=\frac{1}{\sigma_{Q}}$
where $\sigma_{Q}$ are the singular values of the matrix
$Q=I-W.$
Therefore, low-rank suppression occurs whenever $Q$ has at least one asymptotically large singular value,
$\sigma_{Q}\gg 1.$
Since $W$ has an asymptotically large singular value, low-rank suppression occurs because this large singular value is conserved by adding the identity matrix. However, unlike eigenvalues, singular values are not transformed in a simple way by adding the identity matrix, in other words, we do not generally have $\sigma_{Q}=1-\sigma_{W}$. If $W$ were a normal matrix, then we would have $\sigma_{Q}=|1-\sigma_{W}|$, so the emergence of low-rank suppression would have been easier to prove, but this is not in general true of non-normal matrices. The potentially complicated relationship between the singular values of $W$ and those of $Q$ is more critical for the analysis of high-dimensional dynamics in the next section.
The proof of Claim 1 also provides insight into the input and response directions that are suppressed in low-rank suppression. Note that ${\bf{x}}_{eff}={\bf{u}}_{k}+o(1)$ and ${\bf{z}}\propto{\bf{v}}_{k}$, which implies that inputs approximately aligned to the column space of $U$ (equivalently, the column space of $W_{0}$) give rise to suppressed responses aligned to the column space of $V$ (equivalently, the row space of $W_{0}$).
Finally, it is worth discussing our use of ${\bf{x}}_{eff}$ instead of ${\bf{x}}$ in the definition of low-rank suppression. For our simulations ${\bf{x}}={\bf{x}}_{eff}$ so there is no issue. However, one could consider a model for which $J_{{\bf{x}}}$ is full rank, but amplifies some directions more than others. For example, consider the model in Eq. (11) with a $W_{x}$ chosen to amplify inputs in every direction parallel to the column space of $W$, i.e.., if $\|W_{x}{\bf{u}}_{k}\|\gg 1$ for $k=1,\ldots,r$ whereas $\|W_{x}{\bf{u}}_{rand}\|=\mathcal{O}(1)$ for random directions. In this very specific case, the amplification by $W_{x}$ would compete with the suppression by $A=[I-W]^{-1}$, so the raw input, ${\bf{x}}$, would need to be smaller in the direction of ${\bf{u}}_{k}$ to obtain an ${\bf{x}}_{eff}$ that is $\mathcal{O}(1)$ in that direction. In this case, the suppression by recurrent dynamics would compete with the amplification by a feedforward read-in. Importantly, the recurrent dynamics would still be suppressive to effective inputs aligned to ${\bf{u}}_{k}$. We focus on effective inputs in our definition of low-rank suppression to avoid such an issue, but it would only arise in a very specific situation where $J_{{\bf{x}}}$ amplifies exactly the directions that $A=[I-W]^{-1}$ suppresses.
Above, we considered slow or static perturbations, ${\bf{x}}(t)$. Specifically, we assumed that perturbations were constant in time or changed more slowly than the intrinsic dynamics of the network (quantified by $\tau$ in our models). This assumption applies to all of the examples used in the text, but in some applications, perturbations are faster than intrinsic network dynamics. We now consider transient perturbations, ${\bf{x}}(t)$, with arbitrary timescales. We can no longer rely on the quasi-static approximation from Eq. (13) in this case. Instead, we take the Laplace transform in Eq. (9) to obtain
$\hat{{\bf{z}}}=\hat{H}\hat{{\bf{z}}}+\hat{B}\hat{{\bf{x}}}$ (16)
where
$\hat{H}(s)$ $=(1+\tau s)^{-1}\left[J_{{\bf{z}}}+I\right]=(1+\tau s)^{-1}W$ (17)
$\hat{B}(s)$ $=(1+\tau s)^{-1}J_{{\bf{x}}}$
[…]
The $N\times N$ matrix $\hat{H}(s)$ can be interpreted as a measure of the effective connectivity of the linearized network at mode $s$, i.e.,
$W=\hat{H}(s)$
because each entry $\hat{H}_{jk}(s)$, represents the linearized response of ${\bf{z}}_{j}(t)$ to fluctuations in ${\bf{z}}_{k}(t)$ at Laplace mode $s$. Solving the implicit Eq. (16) for $\hat{{\bf{z}}}$ gives the response to arbitrary inputs in the Laplace domain,
$\hat{{\bf{z}}}=[I-W]^{-1}\hat{B}\hat{{\bf{x}}}.$ (18)
Since $\hat{B}$ quantifies feedforward, read-in dynamics, we can interpret $\hat{B}\hat{{\bf{x}}}$ as effective input in the Laplace domain, ${\bf{\hat{x}}}_{eff}=\hat{B}\hat{{\bf{x}}}$. This is analogous to our treatment of $J_{{\bf{x}}}{\bf{x}}$ above.
Eq. (18) is identical to Eq. (13) except that Eq. (18) quantifies interactions at individual Laplace modes, $s$. Hence, Claim 1 applies to situations in which ${\bf{x}}(t)$ is transient and fast, so long as the effective connectivity, $W$, and effective input, ${\bf{x}}_{eff}$ are interpreted in the Laplace domain ($W=H(s)$ and ${\bf{x}}_{eff}=\hat{B}(s)\hat{{\bf{x}}}(s)$).
Moreover, the approach of working in the Laplace domain allows us to expand our class of models. In all of the analysis above, we considered dynamics obeying the system of ordinary differential equations in Eq. (7). The dynamics in Eq. (7) capture a large class of models, but are not fully general. Specifically, in Eq. (7), whenever $t_{0}<t_{1}$, the value of ${\bf{z}}(t_{1})$ is fully determined by the value of ${\bf{z}}(t_{0})$ and the values of $x(s)$ for $s\leq t_{1}$. However, some systems like integro-differential equations and non-Markovian stochastic systems have a history dependence that cannot be captured by a system of ordinary differential equations like Eq. (7). In particular, we can weaken our assumptions on the dynamics of ${\bf{z}}(t)$ and assume only that ${\bf{z}}(t)$ is the response of a causal, time-invariant system that satisfies
${\bf{z}}(t)=\mathcal{F}\left(\{{\bf{z}}(s),{\bf{x}}(s)\}_{s<t}\right).$ (19)
In other words, ${\bf{z}}(t)$ is an arbitrary time-translation equivriant function of the history of ${\bf{z}}(s)$ and ${\bf{x}}(s)$. Eq. (19) encompasses an extremely broad class of systems, including those described by Eq. (7).
A first order Volterra expansion gives Eq. (16) as an implicit equation for ${\bf{z}}(t)$. The matrices $\hat{H}$ and $\hat{B}$ are functional derivatives of $\mathcal{F}$, evaluated at the baseline. Under the dynamics in Eq. (9), we obtain the $\hat{H}$ and $\hat{B}$ defined in Eq. (17). Hence, this approach generalizes the results in Claim 1 to fast, transient inputs and to the very general class of dynamics defined by Eq. (19).
S.2.4 Conditions for high-dimensional dynamics in response to stationary, stochastic perturbations.
We now consider conditions for the emergence of high-dimensional dynamics (as in Figure 1d). We assume that ${\bf{x}}(t)$ is a stationary, ergodic stochastic process, as in Figure 1d–g. Define the cross-spectral matrix, $\widetilde{C}^{{\bf{x}}}(f)$, of ${\bf{x}}(t)$ at frequency $f$ as the Fourier transform of the matrix of cross-covariance matrix,
[…]
where $C_{jk}^{{\bf{x}}}(\tau)=\text{cov}({\bf{x}}_{j}(t),{\bf{x}}_{k}(t+\tau))$ is the stationary cross-covariance. When each ${\bf{x}}_{j}(t)$ is i.i.d., then $\widetilde{C}^{{\bf{x}}}(f)=I\widetilde{a}_{x}(f)$ is a multiple of the identity matrix where $\widetilde{a}_{x}(f)$ is the power spectral density of each ${\bf{x}}_{j}(t)$. This is true for all of the examples we consider, but we will not apply this simplification until the end of our calculation.
@@
$\widetilde{C}^{{\bf{z}}}(f)$ $=[\widetilde{r}(f)I-W]^{-1}J_{{\bf{x}}}\widetilde{C}^{{\bf{x}}}(f)J_{{\bf{x}}}^{T}[\widetilde{r}(f)I-W]^{-*}$ (20)
@@
Note that Eq. (20) can be derived from Eqs. (18) and (17) by first taking $s=-2\pi if$ in Eq. (18) to switch from the Laplace domain to the Fourier domain, and then applying the Wiener–Khinchin theorem to write $\widetilde{C}^{{\bf{z}}}(f)=\widetilde{{\bf{z}}}(f)\widetilde{{\bf{z}}}^{*}(f)=\hat{{\bf{z}}}(-2\pi if)\hat{{\bf{z}}}^{*}(-2\pi if)$. Hence, the results derived in this section apply to the general class of dynamics defined in Eq. (19).
@@
[…]
If the timescale of fluctuations in ${\bf{x}}(t)$ are much slower than the timescale, $\tau$, of network dynamics then $\widetilde{C}^{{\bf{x}}}(f)\approx 0$ for $f>\epsilon/\tau$ where $\epsilon\ll 1$. Indeed, we take this to be the definition of the statement that the fluctuations in ${\bf{x}}(t)$ are much slower than $\tau$. Since $\widetilde{r}(f)\approx 1$ whenever $f\tau<\epsilon$, we therefore have that $\widetilde{r}(f)\approx 1$ whenever $\widetilde{C}^{{\bf{x}}}(f)$ is not close to zero. As a result, the only parts of the integrand that contribute to the integral in Eq. (21) are the low frequency components, $f\approx 0$. In this case, we can replace $\widetilde{r}(f)$ with $1$ to obtain
$\overline{C}^{{\bf{z}}}\approx[I-W]^{-1}J_{{\bf{x}}}\overline{C}^{{\bf{x}}}J_{{\bf{x}}}^{T}[I-W]^{-T}.$ (22)
@@
In the models we consider, each ${\bf{x}}_{j}(t)$ is an i.i.d. stochastic process, so $\overline{C}^{{\bf{x}}}=vI$ is a multiple of the identity where $v=\text{var}({\bf{x}}_{j}(t))$ is the stationary covariance of ${\bf{x}}_{j}(t)$. Moreover, in the networks we consider, $J_{{\bf{x}}}=I$ is equal to the identity matrix. Therefore, for our models,
$\overline{C}^{{\bf{z}}}\approx[I-W]^{-1}v$
where $v$ is a scalar.
In more general classes of models in which $\overline{C}^{{\bf{x}}}$ or $J_{{\bf{x}}}$ are not multiples of the identity matrix, dimensionality of the dynamics of ${\bf{z}}(t)$ could be reduced whenever $\overline{C}^{{\bf{x}}}$ or $J_{{\bf{x}}}$ are effectively low-rank. However, this effect would be caused by a low-dimensional perturbation ($\overline{C}^{{\bf{x}}}$ low-dimensional) or a low-rank read-in mechanism ($J_{{\bf{x}}}$ low-dimensional; see Supplementary Figure S.9 and discussion in previous sections), and would not be related to the actual dynamics of interactions within the network. Therefore, our focus on the effective rank of $[I-W]^{-1}$ in place of $[I-W]^{-1}J_{{\bf{x}}}\sqrt{\overline{C}^{{\bf{x}}}}$ is justified even in more general settings.
In conclusion, the dimensionality of the dynamics of ${\bf{z}}(t)$ is determined by the effective rank of the matrix
@@
[…]
where $\sigma_{A,k}$ is the $k$th singular value of $A$ (assuming singular values are sorted in decreasing order) and $v=\text{var}({\bf{x}}_{j}(t))$ is the stationary variance of each ${\bf{x}}_{j}(t)$.
@@
[…]
where $\sigma_{Q,N-k}$ is the $(N-k)$th singular value of $Q$, i.e., the $k$th from the last singular value (assuming singular values are sorted in decreasing order).

### v2

@@ Acknowledgments
S.2 Conditions for high-dimensional dynamics in response to stationary, stochastic perturbations.
We now consider conditions for the emergence of high-dimensional dynamics (as in Figure 1d). We assume that ${\bf{x}}(t)$ is a stationary, ergodic stochastic process, as in Figure 1d–f. Define the cross-spectral matrix, $\widetilde{C}^{{\bf{x}}}(f)$, of ${\bf{x}}(t)$ at frequency $f$ as the Fourier transform of the matrix of cross-covariance matrix,
$\widetilde{C}^{{\bf{x}}}_{jk}(f)=\int_{-\infty}^{\infty}C_{jk}^{{\bf{x}}}(\tau)e^{-2\pi if\tau}d\tau$
where $C_{jk}^{{\bf{x}}}(\tau)=\text{cov}({\bf{x}}_{j}(t),{\bf{x}}_{k}(t+\tau))$ is the stationary cross-covariance. When each ${\bf{x}}_{j}(t)$ is i.i.d., then $\widetilde{C}^{{\bf{x}}}(f)=I\widetilde{a}_{x}(f)$ is a multiple of the identity matrix where $\widetilde{a}_{x}(f)$ is the power spectral density of each ${\bf{x}}_{j}(t)$. The assumption of i.i.d. elements is true for all of the examples we consider, but we will not apply this simplification until the end of our calculation.
@@
$\widetilde{C}^{{\bf{z}}}(f)$ $=[\widetilde{r}(f)I-W]^{-1}\widetilde{C}^{{\bf{x}}}(f)[\widetilde{r}(f)I-W]^{-*}$ (13)
@@
$\overline{C}^{{\bf{z}}}=\int_{-\infty}^{\infty}[\widetilde{r}(f)I-W]^{-1}\widetilde{C}^{{\bf{x}}}(f)[\widetilde{r}(f)I-W]^{-*}df$ (14)
If the timescale of fluctuations in ${\bf{x}}(t)$ are much slower than the timescale, $\tau$, of network dynamics then $\widetilde{C}^{{\bf{x}}}(f)\approx 0$ for $f>\epsilon/\tau$ where $\epsilon\ll 1$. Indeed, we can take this to be the definition of the statement that the fluctuations in ${\bf{x}}(t)$ are much slower than $\tau$. Since $\widetilde{r}(f)\approx 1$ whenever $f\tau<\epsilon$, we therefore have that $\widetilde{r}(f)\approx 1$ whenever $\widetilde{C}^{{\bf{x}}}(f)$ is not close to zero. As a result, the only parts of the integrand that contribute to the integral in Eq. (14) are the low frequency components, $f\approx 0$. In this case, we can replace $\widetilde{r}(f)$ with $1$ to obtain
$\overline{C}^{{\bf{z}}}\approx[I-W]^{-1}\overline{C}^{{\bf{x}}}[I-W]^{-T}.$ (15)
Note that Eq. (15) can also be derived directly from the quasi-steady state approximation ${\bf{z}}(t)\approx[I-W]^{-1}{\bf{x}}(t)$. Our assumption above that $\widetilde{C}^{{\bf{x}}}(f)\approx 0$ for $f>\epsilon/\tau$ is essentially equivalent to the assumption that the quasi-steady state approximation is accurate, but our approach here is more mathematically precise.
@@
In the models we consider, each ${\bf{x}}_{j}(t)$ is an i.i.d. stochastic process, so $\overline{C}^{{\bf{x}}}=vI$ is a multiple of the identity where $v=\text{var}({\bf{x}}_{j}(t))$ is the stationary covariance of ${\bf{x}}_{j}(t)$. Therefore, for our models,
$R=[I-W]^{-1}\sqrt{v}$
and therefore
$\overline{C}^{{\bf{z}}}\approx[I-W]^{-1}[I-W]^{-T}v$
where recall that $v>0$ is a scalar.
In more general classes of models in which $\overline{C}^{{\bf{x}}}$is not a multiple of the identity matrix, the dimensionality of ${\bf{z}}(t)$ might be reduced when $\overline{C}^{{\bf{x}}}$ is effectively low-rank. However, this effect would be caused by a low-dimensional perturbation ($\overline{C}^{{\bf{x}}}$ low-dimensional) (see Supplementary Figure S.2), and would not be related to the intrinsic dynamics of interactions within the network.
In conclusion, when ${\bf{x}}(t)$ is high-dimensional, the dimensionality of the dynamics of ${\bf{z}}(t)$ is determined by the effective rank of the matrix
@@
$\textrm{var. explained by $k$th PC of }{\bf{z}}(t)\approx\sigma^{2}_{A,k}\,v$
where $\sigma_{A,k}$ is the $k$th singular value of $A$ (assuming singular values and principal components are sorted in decreasing order) and $v=\text{var}({\bf{x}}_{j}(t))$ is the stationary variance of each ${\bf{x}}_{j}(t)$.
@@
$\textrm{var. explained by $k$th PC of }{\bf{z}}(t)\approx\frac{v}{\sigma^{2}_{Q,N-k}}$ (16)
where $\sigma_{Q,N-k}$ is the (Nk)(N-k)th singular value of $Q$, i.e., the $k$th from the last singular value.